OpenStudy (anonymous):

The period of a pendulum is given by the formula T=2pi(sqrt(L/g)) where L is the length of the pendulum in feet, g = 32 ft/s2 is the acceleration due to gravity, and T is the length of one period in seconds. (a) If the length of the pendulum is measured to be 3 feet long with a max error of 1/8th of an inch, what is the approximate max error in the period, T? ______________________ (b) Find the relative error in the period, in terms of the relative error in the length. ______________________ (c) Find the relative error in the period, for part a. ______________________

OpenStudy (anonymous):

\[\Delta L=0.125\]\[T(L)=2\pi \sqrt{\frac{L}{g}}=\frac{2\pi}{\sqrt{g}}.\sqrt{L}=k \sqrt{L}\], where \[k=\frac{2\pi}{\sqrt{g}}\], a constant Then approximate small change in T with small change in L is given by \[\Delta T=\frac{dT}{dL} . \Delta L\]